Circular No. 7545

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000fm IN NGC 1612 M. Schwartz, Cottage Grove, OR; and W. D. Li and A. V. Filippenko,

University of California at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova with the 0.5-m Tenagra III automated telescope. The object was discovered and confirmed on unfiltered images taken on Dec. 14.3 (mag ∼ 17.7) and 15.2 UT (mag ∼ 17.0). SN 2000fm is located at α = 4^{h}33^{m}15^{s}.00, δ = −4^{o}10′39′′.5 (equinox 2000.0), which is 24′′.1 east and 19′′.0 south of the nucleus of NGC 1612. An image of the field taken with the Katzman Automatic Imaging Telescope on Dec. 2.4 showed nothing at this position (limiting magnitude ∼ 19.0).

SUPERNOVA 2000fe IN UGC 4870 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics (CfA), report that a spectrum of SN 2000fe (cf. IAUC 7540), obtained by P. Berlind on Dec. 15.37 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova before maximum light. The spectrum consists of an almost featureless blue continuum. Incipient P-Cyg lines of hydrogen (and perhaps helium) are apparent.

S. Benetti, G. Altavilla, E. Cappellaro, S. Desidera, D. Fantinel, E. Giro, L. Lessio, A. Pastorello, C. Pernechele, and M. Turatto, Osservatorio Astronomico di Padova, write: “The fully reduced CCD spectrum (range 330–740 nm, resolution 1.8 nm) of SN 2000fe, obtained on Dec. 19.04 UT with the Asiago 1.82-m telescope (+ AFOSC), is that of a type-II supernova near maximum light. The spectrum exhibits a blue continuum with broad P-Cyg Balmer and He I 587.6-nm lines. The expansion velocity deduced from the Hβ minimum is ∼ 8800 km/s. Hα shows a double broad absorption with minima measured at 629.6 and 644.6 nm, where the first minimum could be identified with the Si II 635.5-nm transition. A similar structure in the Hα absorption has been also noticed in a spectrum of SN 2000N. The galaxy redshift, as measured from emission lines of H II regions located near the supernova position, is 4150 km/s.”

S/2000 S 11 The team led by B. Gladman (cf. IAUC 7512) has reported yet another

new Saturnian satellite, this one discovered by M. Holman, CfA, with the 1.2-m telescope at the F. L. Whipple Observatory on Nov. 9. Complete details are on MPEC 2000-Y13.

2000 December 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green